# AMBER: An Adaptive Multimodal Mask Transformer for Beam Prediction with Missing Modalities

### v1

@@ III-B4 Class-former-aided Modality Alignment
III-C Beam Prediction
[…]
The overall training loss function is defined as
$L=\lambda_{f}L_{f}+\lambda_{c}L_{c}+\lambda_{r}L_{2},$ (38)
IV Simulation Results
IV-A Simulation Settings
TABLE III: comparison of the proposed method with classical baselines based on average Top-3 Accuracy and DBA score
@@ IV-A Simulation Settings
Image+GPS Vision-Position[16] 76.61 65.99 66.26 61.22 62.43 0.8790
@@
CMFE [17] 100.0 69.95 66.35 70.11 76.61 0.9264
AMBER(ours) 100.0 87.71 86.26 87.14 88.25 0.9289
@@
AMBER(ours) 100.0 87.71 86.93 88.50 89.07 0.9306
TABLE IV: Missing Specific Modality
@@
Image 61.0577 86.2500 93.0769
LiDAR 63.8782 88.3654 94.6795
Radar 63.4615 87.9487 94.8397
Beam index 63.6218 89.0705 95.1564
GPS 63.8141 88.8141 94.6154
None 64.1910 89.0705 95.3205
IV-A1 Dataset and training setup
To evaluate the performance of the proposed AMBER framework under realistic mmWave communication conditions, we conduct experiments using the real-world DeepSense 6G dataset [24], which has become a key benchmark for sensing-aided beamforming research. In this dataset, a stationary BS is equipped with an RGB camera, radar, and LiDAR sensor, and serves a mobile UE. The UE is also equipped with a GPS receiver that continuously reports meter-level positional information to the BS.
As summarized in Table I, the dataset include four different scenarios (Scenarios 31-34). They are recorded at street-level in a dynamic urban environment, capturing a variety of real-world dynamics. Scenarios 31 and 32 are recorded during the daytime, while scenarios 33 and 34 are collected at night. All scenarios include vehicles of various sizes and speeds, as well as pedestrian activities, thereby introducing substantial environmental variations. These characteristics make DeepSense 6G a representative and challenging testbed for assessing the robustness and effectiveness of AMBER.
For model training and evaluation, the dataset is randomly divided into 80% training and 20% testing subsets, each corresponding to an independent vehicle pass-by event with synchronized multimodal sensor data and beam index sequences. Ground-truth beam labels are obtained from the mmWave received power vectors at the BS. All experiments are conducted on a single NVIDIA RTX 4090 GPU using the PyTorch framework. The overall simulation and training parameters are listed in Table II. AMBER is trained end-to-end using the loss function $L$ defined in (38). For radar and LiDAR feature extraction, we adopt a pretrained ResNet-18[26]. For the image modality, we initialize the encoder with a pretrained ResNet-34[26], given that the image stream contains richer semantic context and thus play a more critical role in accurate beam prediction (as analyzed in Section IV-B). Leveraging pretrained models can accelerate convergence and reduce overfitting by providing well-regularized initial feature representations.
@@ IV-A2 Evaluation Metrics
TABLE V: Comparison of Model parameter and computational complexity
Method Parameters Flops
TII-Transfuser [18] 66.16M 11.42G
Vision-Position[16] 33.02M 7.14G
BeMamba[22] 16.19M 23.43G
Multimodal-LSTM [19] 49.01M 52.80G
CMFE [17] 74.03M 11.37G
AMBER(ours) 46.52M 47.27G
@@
IV-B Performance Evaluation
@@
Table III presents the performance comparison between the proposed AMBER method and several representative multimodal baselines in terms of Top-3 accuracy and DBA score. The baseline methods used in this study cover diverse modality fusion scenarios. For AMBER, the same network architecture is used across all input settings, with only the modality mask adjusted to reflect the available sensing inputs. In contrast, baseline models are individually designed and trained for each specific modality combination. The results demonstrate that AMBER consistently achieves state-of-the-art performance across all metrics, verifying its effectiveness and its adaptability. Across all scenarios, AMBER achieves an average Top-3 accuracy of 89.07% and a DBA score of 93.06%. Specifically, when all modalities are available, AMBER surpasses Multimodal-LSTM [19] by 6.28% in Top-3 accuracy and 2.86% in DBA score. Under more challenging nightime environments (Scenarios 33 and 34), AMBER achieves 5.54% and 10.18% higher Top-3 accuracy than Multimodal-LSTM, highlighting its strong robustness to degraded lighting and increased sensor noise.
Table V presents a comparison of computational efficiency in terms of parameter size and FLOPs. AMBER achieves a favorable trade-off between performance and complexity, requiring only 46.52M parameters and 47.27G FLOPs, substantially lower than the 66.16M parameters of the transformer-based TII-Transfuser model. These results demonstrate the compactness and efficiency of AMBER’s design.
To further evaluate performance with multiple missing modalities, Fig. 6 presents the results obtained by masking up to three modalities with 50% masking ratio. While both methods achieve higher accuracy under a relaxed Top-$K$ criterion, AMBER consistently outperforms Multimodal-LSTM. As the number of missing modalities increases from 0 to 3, AMBER’s Top-1 accuracy decreases moderately from 64.04% to 60.86%, whereas Multimodal-LSTM’s exhibits a steep decline from 58.49% to 45.03%. The performance degradation rate of Multimodal-LSTM is approximately 4.2 times higher than that of AMBER (13.46% vs. 3.18%), underscoring AMBER’s reliability under incomplete modality conditions. This robustness is primarily attributed to the adaptive masking mechanism, which ensures that the transformer attends only to available modalities and avoids propagating noise from missing ones.
@@
TABLE VI: Ablation study of beam prediction performance.
FullMod Miss2MR05 PosEmb ReweightInd CMA Top-1 Top-3 Top-5
63.46 88.48 95.01
64.02 (+0.56) 88.69 (+0.21) 95.06 (+0.05)
64.11 (+0.65) 88.92 (+0.44) 95.18 (+0.17)
64.19 (+0.73) 89.07 (+0.59) 95.32 (+0.31)
62.19 86.64 94.05
62.32 (+0.13) 86.67 (+0.03) 94.11 (+0.06)
62.47 (+0.28) 86.98 (+0.34) 94.30 (+0.25)
62.89 (+0.70) 87.31 (+0.67) 94.62 (+0.57)
@@
IV-C Ablation Study
In this subsection, an ablation study is conducted to quantify the contributions of three key components in AMBER: the positional embedding, the reweighting indicator, and the cross-modal alignment module. Two settings are considered: (i) all modalities are available and (ii) two randomly selected modalities are missing with a 50% masking ratio. The baseline model is defined as AMBER with all three components removed. The results, summarized in Table VI, reveal similar performance trends across both settings; thus, the following discussion focuses on the full-modality case.
Specifically, introducing positional embeddings improves the model’s ability to capture spatio-temporal dependencies, yielding a 0.56% improvement in Top-1 accuracy. Adding the reweighting indicator further boosts performance by adaptively balancing the relative importance of different modalities, resulting in an additional 0.65% improvement in Top-1 accuracy (from 63.46% to 64.11%). The CMA module deliver the most substantial improvement, raising the Top-1 accuracy to 64.19%, i.e., a 0.73% increase over the baseline. By explicitly aligning multimodal features via contrastive learning, CMA reduces inter-modal discrepancies, enforces semantic consistency, and enhances the model’s ability to exploit complementary information across modalities.
The pooling size determines both the dimension of the modality-specific features (including LiDAR, radar, and camera features) and the length of the transformer input $\bar{\bm{\xi}}$. Its selection reflects a trade-off between model performance and inference efficiency. The impact of the pooling size on accuracy and inference speed is analyzed in Fig. 9. As illustrated in Fig. 9(a), the accuracy initially improves as the pooling size increases but decreases beyond a certain point due to feature redundancy. Larger pooling sizes introduce excessive, repeated information, increasing the risk of overfitting and weakening generalization. Meanwhile, as shown in Fig.9(b), inference time grows quadratically with pooling size, as the number of input tokens scales accordingly. Among all tested configurations, a pooling size of (4, 4) offers the best trade-off, achieving the highest Top-$K$ accuracy while maintaining an efficient inference speed of 26.79 FPS, the feasibility of AMBER for practical deployment.

### v2

@@ IV-A1 Dataset and training setup
In this work, we use four urban street scenarios (Scenarios 31–34), as summarized in Table I, to comprehensively validate the robustness and effectiveness of AMBER under diverse real-world conditions. These scenarios cover different illumination and traffic environments, with Scenarios 31 and 32 collected during the daytime and Scenarios 33 and 34 collected at night. They also include heterogeneous road users, such as vehicles of different sizes and speeds as well as pedestrian activities, thereby introducing substantial environmental dynamics and sensing variability. These characteristics make DeepSense 6G a representative and challenging testbed for assessing the robustness and effectiveness of AMBER.
TABLE III: comparison of the proposed method with classical baselines based on average Top-3 Accuracy (%) and DBA score
@@
Image+GPS Vision-Position[18] 90.38 78.23 83.32 80.14 80.64 0.8203
@@
CMFE [19] 100.0 72.75 71.44 74.73 76.64 0.9264
AMBER(ours) 100.0 87.61 86.16 87.04 88.25 0.9289
@@
AMBER(ours) 100.0 87.48 86.26 88.23 89.07 0.9294
TABLE IV: Impact of Specific Missing Modality on Accuracy (%)
@@
Image 61.06 86.25 93.08
LiDAR 63.88 88.37 94.68
Radar 63.46 87.95 94.84
Beam index 63.78 88.63 95.15
GPS 63.81 88.81 94.62
None 64.15 89.07 95.33
For model training and evaluation, the dataset is randomly divided into 80% training and 20% testing subsets, each corresponding to an independent vehicle pass-by event with synchronized multimodal sensor data and beam index sequences. Ground-truth beam labels are obtained from the mmWave received power vectors at the BS. All experiments are conducted on a single NVIDIA RTX 4090 GPU using the PyTorch framework. The overall simulation and training parameters are listed in Table II. AMBER is trained end-to-end using the loss function $L$ defined in (40). For radar and LiDAR feature extraction, we adopt a pretrained ResNet-18[29]. For the image modality, we initialize the encoder with a pretrained ResNet-34[29], given that the image stream contains richer semantic context and thus play a more critical role in accurate beam prediction (as analyzed in Section IV-B). Leveraging pretrained models can accelerate convergence and reduce overfitting by providing well-regularized initial feature representations. The GPS encoder ${\mathcal{P}}_{G}$ is implemented as a three-layer MLP. Specifically, the normalized GPS input is first projected to 64 a 64-dimensional feature space via a linear layer, followed by LayerNorm and ReLU activation. The feature is then mapped from 64 to 256 dimensions, again followed by LayerNorm and ReLU. Finally, a linear layer projects the feature to the common embedding dimension of 256, followed by a final LayerNorm. The beam encoder ${\mathcal{P}}_{B}$ adopts the same three-layer MLP architecture as ${\mathcal{P}}_{G}$, except that the input dimension of its first linear layer is adjusted to match the dimensionality of the beam features.
@@ IV-A2 Evaluation Metrics
IV-B Performance Evaluation
Table III presents the performance comparison between the proposed AMBER method and several representative multimodal baselines in terms of Top-3 accuracy and DBA score. The baseline methods used in this study cover diverse modality fusion scenarios. For AMBER, the same network architecture is used across all input settings, with only the modality mask adjusted to reflect the available sensing inputs. In contrast, baseline models are individually designed and trained for each specific modality combination. The results demonstrate that AMBER consistently achieves state-of-the-art performance across all metrics, verifying its effectiveness and its adaptability. Across all scenarios, AMBER achieves an average Top-3 accuracy of 89.07% and a DBA score of 92.94%. Specifically, when all modalities are available, AMBER surpasses Multimodal-LSTM [21] by 5.93% in Top-3 accuracy and 2.54% in DBA score. Under more challenging nightime environments (Scenarios 33 and 34), AMBER achieves 4.87% and 9.91% higher Top-3 accuracy than Multimodal-LSTM, highlighting its strong robustness to degraded lighting and increased sensor noise.
To further evaluate performance with multiple missing modalities, Fig. 6 presents the results obtained by masking up to three modalities with 50% masking ratio. While both methods achieve higher accuracy under a relaxed Top-$K$ criterion, AMBER consistently outperforms Multimodal-LSTM. As the number of missing modalities increases from 0 to 3, AMBER’s Top-1 accuracy decreases moderately from 64.15% to 60.86%, whereas Multimodal-LSTM’s exhibits a steep decline from 58.49% to 45.03%. The performance degradation rate of Multimodal-LSTM is approximately 4.2 times higher than that of AMBER (13.46% vs. 3.29%), underscoring AMBER’s reliability under incomplete modality conditions. This robustness is primarily attributed to the adaptive masking mechanism, which ensures that the transformer attends only to available modalities and avoids propagating noise from missing ones.
TABLE V: Comparison of modality combination and computational complexity
Modality Top-1 Flops Latency
BeamIdx+GPS 58.81% 0.077G 7.98ms
BeamIdx+GPS+Radar 59.85% 11.77G 12.33ms
BeamIdx+GPS+LiDAR 60.26% 11.51G 12.05ms
BeamIdx+GPS+Image 63.42% 24.14G 15.58ms
BeamIdx+GPS+Image+Radar 63.88% 35.83G 19.03ms
Full Modality 64.15% 47.27G 22.64ms
@@ IV-B Performance Evaluation
We further report the inference FLOPs and inference latency under different modality combinations in Table V. The full-modality configuration achieves the highest Top-1 accuracy of 64.15%, but also incurs the largest computational cost (47.27G FLOPs) and latency (22.64 ms). In contrast, lightweight configurations such as BeamIdx+GPS require negligible computation (0.077G) and achieve the lowest latency (7.98ms), at the cost of reduced prediction accuracy. Among the evaluated settings, BeamIdx+GPS+Image provides a favorable accuracy-efficiency trade-off, achieving 63.42% Top-1 accuracy while significantly reducing computational (24.14G FLOPs) and latency (15.58ms) compared to the full-modality setting. These results suggest that the proposed framework is not restricted to a fixed deployment mode, rather, it supports flexible modality combinations to accommodate diverse vehicular hardware constraints and real-time requirements. It is also observed that, although the theoretical FLOPs increase substantially, the measured inference latency grows more moderately. This discrepancy arises because practical GPU latency is influenced not only by arithmetic complexity, but also by factors such as memory access, kernel launch overhead, and parallel execution efficiency. Consequently, while FLOPs and latency exhibit consistent trends, they are not strictly proportional.
@@
IV-C Ablation Study
In this subsection, we conduct an ablation study to quantify the contributions of several key components in AMBER, including the mask module, positional embedding, reweighting indicator, and cross-modal alignment module. Two incomplete-modality settings are considered: (i) Miss1MR05, where one randomly selected modality is missing with a masking ratio of 50%; and (ii) Miss2MR05, where two randomly selected modalities are missing with the same masking ratio. For clarity, we collectively refer to the positional embedding, reweighting indicator, and CMA as robustness enhancement components (RCE). Although these components operate at different stages of the network, they are grouped together due to their shared objective of improving model robustness.
The ablation results are summarized in Table VI. Under the Miss1MR05 setting, introducing the mask module improves the Top-3 accuracy from 84.35% to 87.15%, corresponding to a gain of 2.80%. Incorporating RCE further increases the Top-3 accuracy to 87.84%, yielding an additional improvement of 0.69%. Under the more challenging Miss2MR05 setting, the mask module provides an even larger Top-3 improvement of 3.83%. These results indicate that the mask module contributes the dominant performance gain and plays a key role in handling missing modalities, particularly as the degree of modality absence increases. In comparison, RCE serves as a complementary mechanism that further enhances feature representation and improves the robustness of multimodal fusion.
TABLE VI: Ablation study of beam prediction accuracy (%).
Modality Mask RCE Top-1 Top-3 Top-5
60.38 84.35 93.20
Miss1MR05 62.44 87.15 93.97
63.13 87.84 94.86
57.11 82.81 91.25
Miss2MR05 62.19 86.64 94.05
62.89 87.31 94.62
The pooling size determines both the dimension of the modality-specific features (including LiDAR, radar, and camera features) and the length of the transformer input $\bar{\bm{\xi}}$. Its selection reflects a trade-off between model performance and inference efficiency. The impact of the pooling size on accuracy and inference speed is analyzed in Fig. 9. The accuracy initially improves as the pooling size increases but decreases beyond a certain point due to feature redundancy. Larger pooling sizes introduce excessive, repeated information, increasing the risk of overfitting and weakening generalization. Meanwhile, as shown in Fig.9(b), the overall runtime grows quadratically with approximately the pooling size, since the number of input tokens scales accordingly. Among all evaluated configurations, a pooling size of (4, 4) achieves the best trade-off, delivering the highest Top-$K$ accuracy while maintaining an end-to-end validation throughput of 26.79 FPS. This throughput is measured over the complete validation pipeline, including data loading, CPU-to-GPU transfer, forward inference, and metric computation.